OpenStudy (davidusa):

@Whitemonsterbunny17 I need serious math help with asymptotes of functions and stuff.

OpenStudy (davidusa):

I can't stop thinking about this math problem its driving me crazy!

OpenStudy (anonymous):

me neither

OpenStudy (davidusa):

Lol. @satellite73 My question is about a specific function/graph that changes so much with just an extra constant of 10.

OpenStudy (davidusa):

If you go to https://www.desmos.com/calculator and graph these 2 functions, you will see my problem! \[\frac{ 2x+1 }{ x+3} and \frac{ 2x+11 }{ x+3 }\] How does adding a constant of only 10 change the function so much? @Whitemonsterbunny17

OpenStudy (davidusa):

Here is the picture @Whitemonsterbunny17 @satellite73

OpenStudy (davidusa):

Please help me @mathslover @phi @the_fizicx99 @ipwnbunnies

OpenStudy (anonymous):

@madyhughes

OpenStudy (anonymous):

i was tagging myself?

OpenStudy (davidusa):

ok

OpenStudy (anonymous):

Sorry i dont quite understand what you're supposed to be doing.

OpenStudy (anonymous):

this is certainly not the graph of a function is it?

OpenStudy (aum):

Try 2x+6 in the numerator!

OpenStudy (aum):

Let f(x) = (2x + k) / (x + 3). Domain of f(x) excludes x = -3 where there is a vertical asymptote. f'(x) = { (x+3)(2) - (2x + k)(1) } / (x+3)^2 = (2x + 6 - 2x - k) / (x+3)^2 f'(x) = (6-k) / (x+3)^2 The denominator is ALWAYS positive in the domain of the function. (that is, x = -3 is excluded) The numerator is positive IF k < 6 and negative if k > 6. Therefore, f(x) is an increasing function if k < 6 and f(x) is a decreasing function if k > 6. f(x) is undefined at x = -3 In your first function you had k = 1 and therefore it is always an increasing function. In your second function you had k = 11 and therefore it is always an decreasing function. Hence these two graph differently.

OpenStudy (davidusa):

Never mind, I understand now. The graph is of 2 functions, Each one a different color.

OpenStudy (davidusa):

Aum wins the medal for most rigorous explaination!

OpenStudy (davidusa):

Yay for aum!!! :D

OpenStudy (aum):

Thanks! :D